# Stability of time-periodic solutions to the Navier-Stokes-Fourier system

### v1

@@ 3 Existence and stability of time-periodic solution
This section is devoted to the proof of Theorem 1.1. In order to prove Theorem 1.1, we shall investigate the initial value problem of (1):
@@ 3.1 A priori estimate
The following lemma will be used to derive the time-weighted estimate. The proof of Lemma 3.4 below is the same as in [MR4803477, Proposition 5.2].
@@ 3.2 Local existence
We now turn to prove Theorem 1.1. The following construction of time-periodic solutions is inspired by Valli [MR753158], which concerns a time-periodic problem for the isentropic compressible Navier-Stokes equation in a bounded domain.
@@
Taking limit as $n\to\infty$ on both sides of (44) yields $U_{T}(T)=U^{*}_{\infty}$. By the local in time uniqueness in Proposition 3.9, we conclude that $(\rho_{T},v_{T},\theta_{T})$ is the desired time-periodic solution. The decay estimate (((ii))) follows immediately from Lemma 2.1 (iii), Lemma 2.2 and Proposition 3.2. ∎

### v2

@@ 3 Existence and stability of time-periodic solution
This section is devoted to the proof of Theorem 1.1. We shall investigate the initial value problem of (1):
@@ 3.1 A priori estimate
The following lemma will be used to derive the time-weighted estimate. The proof of Lemma 3.4 below is the same as in [3, Proposition 5.2].
@@ 3.2 Local existence
We now turn to prove Theorem 1.1. The following construction of time-periodic solutions is inspired by Valli [16], which concerns a time-periodic problem for the isentropic compressible Navier-Stokes equation in a bounded domain.
@@
Taking limit as $n\to\infty$ on both sides of (45) yields $U_{T}(T)=U^{*}_{\infty}$. By the local in time uniqueness in Proposition 3.9, we conclude that $(\rho_{T},v_{T},\theta_{T})$ is the desired time-periodic solution. The decay estimate (((ii))) in Theorem 1.1 follows immediately from Lemma 2.2 (iii), Lemma 2.3 and Proposition 3.2. ∎